OpenStudy (anonymous):

difference in simplest form? (n^2+3n+2/n^2+6n+8) - (2n/n+4)

OpenStudy (anonymous):

@evanellis8888

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

HELP! please

OpenStudy (gorv):

\[\frac{ (n+2)(n+1) }{ (n+4)(n+2) }- \frac{ 2n }{ n+4 }\]

OpenStudy (gorv):

\[\frac{ n+1 }{ n+4 }-\frac{ 2n }{ n+4 }=\frac{ n+1-2n }{ n+4 }=\frac{ -n+1 }{ n+4 }\]

OpenStudy (anonymous):

uhhhh so thats it?

OpenStudy (anonymous):

hello?

OpenStudy (gorv):

yeah......sorry for late

OpenStudy (gorv):

today problem with my net

OpenStudy (anonymous):

ok so its -n+1/n+4?

OpenStudy (gorv):

yeah

OpenStudy (anonymous):

XD thanks